OpenStudy (anonymous):

5square root 32(x-8)^5=

OpenStudy (hba):

\[5\sqrt{32(x-8)^5}\] Is that your question ?

OpenStudy (anonymous):

yes

OpenStudy (hba):

Now see this, \[5\sqrt{2 \times 2 \times 2 \times 2 \times 2 (x-8)^5}\] \[5\sqrt{2^2 \times 2^2 \times (x-8)^5}\] \[5[2^2 \times 2^2 \times (x-8)^5]^{1/2}\] I hope you can simplify it further.

OpenStudy (anonymous):

no c an you help me

OpenStudy (hba):

\[5(2^2)^{1/2} \times (2^2)^{1/2} \times [(x-8)^5]^{1/2}\] Now you can simplify it further :)

OpenStudy (anonymous):

no I do not know how

OpenStudy (hba):

Just cancel the powers See \[[a^{1/2}]^2=a\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thenwhat do id o

OpenStudy (hba):

Nothing that is your answer.

OpenStudy (anonymous):

what is thean swer

OpenStudy (anonymous):

a isthe a nswer

OpenStudy (anonymous):

so a is the answer

OpenStudy (anonymous):

thanks